# Cascaded Cross-Modal Transformer for Audio-Textual Classification

### v1

@@ 1 Introduction
Figure 1: Our multimodal pipeline for speech classification. Audio is the original input modality, from which we extract tokens with the Wav2Vec2.0 [1] network that processes audio data in the time domain. To generate the first extra (text) modality, we initially employ an ASR model to transcribe each audio sample into text. We assume that the spoken language is known a priori, and, in the depicted example, the audio is in French [Fr]. The French text is given as input to the CamemBERT [2] language model. The next step is to generate the second text modality, which relies on the FLAN [3] model to translate the French transcripts into English. For the English language modality [En], the text is processed by the BERT model [4]. The audio tokens returned by Wav2Vec2.0 and the text tokens produced by CamemBERT and BERT are further fed into our cascaded cross-modal transformer (CCMT). The final class token provided by CCMT is fed into the classification head. Our framework operates similarly if the original spoken language is English.
@@ 2.3 Multimodal classification
Multimodal deep learning has attracted significant attention in recent years as it enables effective modeling and analysis of complex data from multiple modalities. While research has extensively explored the fusion of audio-visual data [36, 7, 37, 12] and audio-visual-textual data [38, 5], with notable advancements in fusion techniques, the combination of audio and text has received comparatively less attention [9, 39, 40, 17]. Singh et al. [39] proposed a hierarchical approach for multimodal speech emotion recognition, combining 33 audio features with textual features extracted from large language models. In [40], the authors introduced AudiBERT, a model that integrates pretrained audio and text representation models along with a dual self-attention mechanism. In contrast to existing methods, which only aggregate information from fine-tuned audio and text models, we propose to construct additional text modalities through automatic speech recognition and neural machine translation. By incorporating multilingual text branches, our model can capture nuances present in different languages, providing a more comprehensive representation of the original audio data.
One key aspect in multimodal learning is the fusion methodology. A wide variety of frameworks have been proposed in the literature, including approaches based on early, intermediate, and late fusion [41, 42, 43, 44, 45, 46, 47]. For instance, Li et al. [45] introduced an efficient and flexible multimodal fusion method specifically designed for fusing unimodal pretrained transformers. Shvetsova et al. [48] proposed a modality-agnostic fusion transformer that learns to exchange information among multiple modalities, such as video, audio, and text, and integrates them into a fused representation in a joint multimodal embedding space. In a different study, Lee et al. [49] presented a cross-modality attention transformer and a multimodal fusion transformer, which identify correlations between data modalities and perform feature fusion. In a similar fashion, Liu et al. [50] proposed a cross-scale cascaded multimodal fusion transformer to facilitate interaction and fusion among modalities with multi-scale features. In contrast to the aforementioned works, our approach introduces a cascaded cross-attention transformer that incorporates attention across distinct languages, followed by cross-attention between the resulting linguistic features and the speech features. This design allows us to benefit from linguistic representations in different languages and effectively integrate them with the audio modality.
@@ 3.1 Audio branch
We utilize the pretrained Wav2Vec2.0 model [1] to learn representative tokens for the audio modality. The raw audio data is partitioned into $k_{a}\in\mathbb{N^{+}}$ chunks, where $k_{a}$ depends on the input length and differs from one sample to another. The model processes the initial tokens, leading to a meaningful audio representation. This process involves several stages, including preliminary convolutional layers for local feature extraction, subsequent transformer layers for global contextual information extraction, and a quantization step for efficient representation. The output of the Wav2Vec2.0 neural network has the same number of tokens (as the input), denoted as $k_{a}$, which represent the acoustic features of the audio modality. The resulting tokens encapsulate features about the audio signals, including pitch, frequency, and intensity. We highlight that encoders for different modalities produce varying numbers of output tokens. However, our CCMT model requires an equal number of tokens for each modality. To meet the data uniformity requirement imposed by CCMT, we randomly sample a fixed number of tokens, denoted as $k\in\mathbb{N^{+}}$, where $k\leq k_{a}$. The selected tokens are then fed into the audio branch of the CCMT model. The tokens are further combined with tokens from the text modalities to facilitate the multimodal analysis. By fine-tuning the Wav2Vec2.0 model, our framework benefits from a rich representation of the audio modality, enabling the CCMT model to effectively capture and integrate both linguistic and acoustic information.
@@ 4.1 Data sets
HarperValleyBank. The HarperValleyBank (HVB) data set [53] is composed of simple consumer-bank interactions, comprising roughly 23 hours of audio from 1,446 conversations between 59 different speakers. The conversations are simulated and follow realistic templates with controlled complexity and a limited vocabulary size of 700 unique words. We randomly split the data based on speaker identifiers into training (70%), validation (15%) and test (15%), thus avoiding speaker leakage across subsets. We tackle two multi-way classification tasks on HVB: action recognition (classification into 16 classes) and intent recognition (classification into 8 classes).
@@ 4.6 Implementation details
Table 1: Results on the ComParE RSC development set, featuring various audio-based models. The models are either trained from scratch or fine-tuned. The architectures marked with *** use pretrained weights. We present the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
@@
$*{}^{*}$ [1] Time domain $\textbf{71.64}\pm 0.16$ $\textbf{58.12}\pm 0.16$
@@ 4.7.1 Preliminary results for the audio modality
Table 2: Results on the ComParE RSC development set using various language models fine-tuned on French transcripts generated by Wav2Vec2.0 [1] and Whisper [10] ASR models. Whisper S+M+L denotes our augmentation technique, incorporating transcripts from all three ASR models (small, medium and large). We present the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
@@ 4.7.2 Preliminary results for the French text modality
We outline the results obtained for the French text transcripts with various NLP models in Table 2. We explore various ASR models, including Wav2Vec2.0 [1] and three versions of the Whisper architecture [10] (small, medium, and large). When using the Wav2Vec2.0 ASR model, the CamemBERT model [2] achieves the best performance, with a Request UAR of $77.45\%$ and a Complaint UAR of $60.15\%$. For the subsequent experiments, we thus select the CamemBERT model in detriment of the LSTM and FlauBERT language models.
@@ 4.7.4 Comparative results for multimodal methods
Table 6: Ablation study on the ComParE RSC development set, considering different configurations of the CCMT architecture trained on all three modalities. We vary the dimension of the projection layer, the number of transformer blocks (depth), the number of cross-attention heads, and the token dimension in each head. We report the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
Projection Depth Heads Head UAR
dimension dimension Request Complaint
256 6 8 512 $81.14\pm 0.11$ $62.98\pm 0.13$
512 6 8 512 $81.45\pm 0.10$ $63.21\pm 0.11$
- 6 8 512 $82.97\pm 0.08$ $66.18\pm 0.09$
- 7 8 512 $83.28\pm 0.08$ $66.11\pm 0.07$
- 8 8 512 $\textbf{83.31}\pm 0.08$ $\textbf{66.64}\pm 0.08$
- 9 8 512 $83.22\pm 0.08$ $65.99\pm 0.08$
- 8 7 512 $83.11\pm 0.07$ $66.65\pm 0.08$
- 8 9 512 $83.19\pm 0.08$ $66.49\pm 0.07$
- 8 8 256 $83.22\pm 0.08$ $66.02\pm 0.08$
- 8 8 1024 $83.01\pm 0.08$ $65.31\pm 0.09$
@@ 4.7.6 Ablation study
Furthermore, we explore the impact of varying the depth, the number of attention heads, and the token dimension in each head of the CCMT architecture. Interestingly, we find that these variations do not lead to significant performance changes. The best results are obtained with a depth of 8 blocks ($L_{1}=L_{2}=8$), each with 8 attention heads and a dimension of 512 for the tokens. Overall, the ablation study emphasizes the importance of removing the input projection in the CCMT model, while indicating that other variations have a relatively minor performance impact.
@@ 4.8 Results on the Speech Commands v2 data set
In Table 7, we present the results of the proposed framework versus the results of several state-of-the-art methods on the Speech Commands v2 data set. We include results for multiple modalities, such as audio-only, text-only and distinct combinations of multimodal inputs. To the best of our knowledge, there is no previous work that uses audio-textual data on this data set. Therefore, we can only compare with audio-based state-of-the-art methods proposed in recent studies [57, 21, 22]. We observe that the Wav2Vec2.0 model surpasses two recent methods [57, 21], while the text-only BERT achieves lower results. However, our audio-textual model outperforms all state-of-the-art methods, regardless of the inclusion of the French text modality. When we combine the English and French text modalities, there is no statistically significant gain over using the English modality alone. The main reason behind this fact is that the data set mostly consists of single-word samples, and changing the language by translation brings no significant contribution. Our best accuracy of $98.86\%$ is obtained by combining audio and text data in both English and French. With this score, CCMT sets a new accuracy record on the Speech Commands v2 data set.

### v2

@@ 1 Introduction
Figure 1: Our multimodal pipeline for speech classification. Audio is the original input modality, from which we extract tokens with the Wav2Vec2.0 [1] network that processes audio data in the time domain. To generate the first extra (text) modality, we initially employ an ASR model to transcribe each audio sample into text. We assume that the spoken language is known a priori, and, in the depicted example, the audio is in French [Fr]. The French text is given as input to the CamemBERT [2] language model. The next step is to generate the second text modality, which relies on the FLAN [3] model to translate the French transcripts into English. For the English language modality [En], the text is processed by the BERT model [4]. The audio tokens returned by Wav2Vec2.0 and the text tokens produced by CamemBERT and BERT are further fed into our cascaded cross-modal transformer (CCMT). The final class token provided by CCMT is fed into the classification head. Our framework operates similarly if the original spoken language is English. The frozen models are marked with a snowflake.
@@ 2.3 Multimodal classification
Multimodal deep learning has attracted significant attention in recent years as it enables effective modeling and analysis of complex data from multiple modalities. While research has extensively explored the fusion of audio-visual data [36, 7, 37, 12] and audio-visual-textual data [38, 5], with notable advancements in fusion techniques, the combination of audio and text has received comparatively less attention [9, 39, 40, 17]. Singh et al. [39] proposed a hierarchical approach for multimodal speech emotion recognition, combining 33 audio features with textual features extracted from large language models. In [40], the authors introduced AudioBERT, a model that integrates pretrained audio and text representation models along with a dual self-attention mechanism. In contrast to existing methods, which only aggregate information from fine-tuned audio and text models, we propose to construct additional text modalities through automatic speech recognition and neural machine translation. By incorporating multilingual text branches, our model can capture nuances present in different languages, providing a more comprehensive representation of the original audio data.
One key aspect in multimodal learning is the fusion methodology. A wide variety of frameworks have been proposed in the literature, including approaches based on early, intermediate, and late fusion [41, 42, 43, 44, 45, 46, 47, 48]. For instance, Li et al. [45] introduced an efficient and flexible multimodal fusion method specifically designed for fusing unimodal pretrained transformers. Shvetsova et al. [49] proposed a modality-agnostic fusion transformer that learns to exchange information among multiple modalities, such as video, audio, and text, and integrates them into a fused representation in a joint multimodal embedding space. In a different study, Lee et al. [50] presented a cross-modality attention transformer and a multimodal fusion transformer, which identify correlations between data modalities and perform feature fusion. In a similar fashion, Liu et al. [51] proposed a cross-scale cascaded multimodal fusion transformer to facilitate interaction and fusion among modalities with multi-scale features. Singh et al. [48] proposed a multimodal fusion framework for emotion recognition by using audio and text features captured by transcribing raw audio files. Similarly, Braunschweiler et al. [52] proposed a multimodal emotion recognition system based on speech and transcribed text, which attained superior results compared to prior studies. Moreover, they showed that the multimodal audio-text fusion leads to better results. In a similar fashion, Sharma et al. [53] proposed a multimodal framework based on video, audio and transcribed text for real-time emotional health detection. In contrast to the aforementioned works, our approach introduces a cascaded cross-attention transformer that incorporates attention across distinct languages, followed by cross-attention between the resulting linguistic features and the speech features. This design allows us to benefit from linguistic representations in different languages and effectively integrate them with the audio modality.
@@ 3.1 Audio branch
We utilize the pretrained Wav2Vec2.0 model [1] to learn representative tokens for the audio modality. The raw audio data is partitioned into $k_{a}\in\mathbb{N^{+}}$ chunks, where $k_{a}$ depends on the input length and differs from one sample to another. The model processes the initial tokens, leading to a meaningful audio representation. This process involves several stages, including preliminary convolutional layers for local feature extraction and transformer layers for global contextual information extraction. The output of the Wav2Vec2.0 neural network has the same number of tokens (as the input), denoted as $k_{a}$, which represent the acoustic features of the audio modality. The resulting tokens encapsulate features about the audio signals, including pitch, frequency, and intensity. We highlight that encoders for different modalities produce varying numbers of output tokens. However, our CCMT model requires an equal number of tokens for each modality. To meet the data uniformity requirement imposed by CCMT, we randomly sample a fixed number of tokens, denoted as $k\in\mathbb{N^{+}}$, where $k\leq k_{a}$. The selected tokens are then fed into the audio branch of the CCMT model. The tokens are further combined with tokens from the text modalities to facilitate the multimodal analysis. By fine-tuning the Wav2Vec2.0 model, our framework benefits from a rich representation of the audio modality, enabling the CCMT model to effectively capture and integrate both linguistic and acoustic information.
@@ 4.1 Data sets
HarperValleyBank. The HarperValleyBank (HVB) data set [56] is composed of simple consumer-bank interactions, comprising roughly 23 hours of audio from 1,446 conversations between 59 different speakers. The conversations are simulated and follow realistic templates with controlled complexity and a limited vocabulary size of 700 unique words. We randomly split the data based on speaker identifiers into training (70%), validation (15%) and test (15%), thus avoiding speaker leakage across subsets. We tackle two multi-way classification tasks on HVB: action recognition (classification into 16 classes) and intent recognition (classification into 8 classes). Wu et al. [56] reported the F1 score for the action recognition task, where the class distribution is highly imbalanced, and the accuracy for the intent recognition task, where the class distribution is almost uniform. Therefore, we report the same performance measures.
@@ 4.6 Implementation details
Table 1: Results on the ComParE RSC development set, featuring various audio-based models. The models are either trained from scratch or fine-tuned. The architectures marked with * use pretrained weights. We present the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
@@
Wav2Vec2.0 [1] Time domain $\textbf{71.64}\pm 0.16$ $\textbf{58.12}\pm 0.16$
@@ 4.7.1 Preliminary results for the audio modality
Table 2: Results on the ComParE RSC development set using various language models fine-tuned on French transcripts generated by Wav2Vec2.0 [1] and Whisper [10] ASR models. Whisper S+M+L denotes our augmentation technique, incorporating transcripts from all three ASR models (small, medium and large). The Wav2Vec2.0 and Whisper models are frozen. We present the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
@@ 4.7.2 Preliminary results for the French text modality
We outline the results obtained for the French text transcripts with various NLP models in Table 2. We explore various ASR models, including Wav2Vec2.0 (large) [^2] https://huggingface.co/facebook/wav2vec2-large-xlsr-53-french [1] and three versions of the Whisper architecture [10] (small, medium, and large). When using the Wav2Vec2.0 ASR model, the CamemBERT model [2] achieves the best performance, with a Request UAR of $77.45\%$ and a Complaint UAR of $60.15\%$. For the subsequent experiments, we thus select the CamemBERT model in detriment of the LSTM and FlauBERT language models.
@@ 4.7.4 Comparative results for multimodal methods
Table 6: Ablation study on the ComParE RSC development set, considering different configurations of the CCMT architecture trained on all three modalities. We vary the dimension of the projection layer, the number of transformer blocks (depth), the number of cross-attention heads, the token dimension in each head, and the token sampling strategy. We report the mean UAR (in percentages) and the standard deviation over three runs, using bold font to indicate the best score on each task.
Projection Depth Heads Head Random UAR
dimension dimension selection Request Complaint
256 6 8 512 $\checkmark$ $81.14\pm 0.11$ $62.98\pm 0.13$
512 6 8 512 $\checkmark$ $81.45\pm 0.10$ $63.21\pm 0.11$
- 6 8 512 $\checkmark$ $82.97\pm 0.08$ $66.18\pm 0.09$
- 7 8 512 $\checkmark$ $83.28\pm 0.08$ $66.11\pm 0.07$
- 8 8 512 $\checkmark$ $\textbf{83.31}\pm 0.08$ $\textbf{66.64}\pm 0.08$
- 9 8 512 $\checkmark$ $83.22\pm 0.08$ $65.99\pm 0.08$
- 8 7 512 $\checkmark$ $83.11\pm 0.07$ $66.63\pm 0.08$
- 8 9 512 $\checkmark$ $83.19\pm 0.08$ $66.49\pm 0.07$
- 8 8 256 $\checkmark$ $83.22\pm 0.08$ $66.02\pm 0.08$
- 8 8 1024 $\checkmark$ $83.01\pm 0.08$ $65.31\pm 0.09$
- 8 8 512 $82.51\pm 0.11$ $66.13\pm 0.09$
@@ 4.7.6 Ablation study
Furthermore, we explore the impact of varying the depth, the number of attention heads, and the token dimension in each head of the CCMT architecture. Interestingly, we find that these variations do not lead to significant performance changes. The best results are obtained with a depth of 8 blocks ($L_{1}=L_{2}=8$), each with 8 attention heads and a dimension of 512 for the tokens.
In the proposed CCMT architecture, after the multi-head attention module, we sum the result with the initial tokens, which generated the keys and values. Considering that the output of the multi-head attention layer has the same length as the query, we cannot sum it with the initial tokens (which generated the keys and values), without having uniformity across the length dimension. In order to perform the ablation study without the random selection strategy, we need to modify the CCMT architecture to be able to handle distinct token dimensions across all data modalities. More exactly, we change the summation operation after both multi-head attention layers inside CCMT, such that the summation is performed using the initial query tokens, instead of the corresponding key and value tokens. We observe that the results are slightly worse, highlighting the efficiency of our architecture based on random token selection.
Overall, the ablation study emphasizes the importance of removing the input projection in the CCMT model, while indicating that other variations have a relatively minor performance impact.
@@ 4.8 Results on the Speech Commands v2 data set
In Table 7, we present the results of the proposed framework versus the results of several state-of-the-art methods on the Speech Commands v2 data set. We include results for multiple modalities, such as audio-only, text-only and distinct combinations of multimodal inputs. To the best of our knowledge, there is no previous work that uses audio-textual data on this data set. Therefore, we can only compare with audio-based state-of-the-art methods proposed in recent studies [60, 21, 22]. We observe that the Wav2Vec2.0 model surpasses two recent methods [60, 21], while the text-only BERT achieves lower results. However, our audio-textual model outperforms all state-of-the-art methods, regardless of the inclusion of the French text modality. When we combine the English and French text modalities, there is no statistically significant gain over using the English modality alone. The main reason behind this fact is that the data set mostly consists of single-word samples, and changing the language by translation brings no significant contribution. Moreover, the ASR models fail in edge case scenarios, e.g. trimmed or muffled words, thus affecting the overall accuracy. To verify this statement, we compute the accuracy by detecting if the words are present in the transcript. The resulting accuracy rate is $94.98\%$. Our best accuracy of $98.86\%$ is obtained by combining audio and text data in both English and French. With this score, CCMT sets a new accuracy record on the Speech Commands v2 data set.